Evaluate the ∫e^z (sq rt of 1+e^z) dz


1
Expert's answer
2021-02-24T12:52:37-0500

Solution.

"\int e^z\sqrt{1+e^z}dz."

Make a replacement:

"t=\sqrt{1+e^z}, t\geq0," from here "t^2=1+e^z, \text{then }
2tdt=e^zdz."

We will have

"\int t\cdot 2tdt=\int 2t^2dt=\newline=
2\frac{t^3}{3}+C=
\frac{2}{3}(1+e^z)^{\frac{3}{2}}+C."

Answer.

"\frac{2}{3}(1+e^z)^{\frac{3}{2}}+C."